Find f(x) and g(x) so that the function can be described as y = f(g(x)).

y = eight divided by x squared + 2
1
Expert's answer
2014-09-08T14:01:08-0400
This problem has many solutions. One of examples is the following: f(x)=x+2, g(x)=8/x2 .